OpenStudy (anonymous):

Tristan says that solutions to radical equations are extraneous. Hadley says that solutions to radical equations are non-extraneous. Find an equation where Tristan is correct and a second equation where Hadley is correct. Using complete sentences, explain each step when solving to justify your examples.

OpenStudy (anonymous):

Does anyone know what extraneous is?

OpenStudy (anonymous):

An extraneous solution is a root of a transformed equation that is not a root of the original equation because it was excluded from the domain of the original equation.

OpenStudy (anonymous):

Does anyone know what extraneous is?

OpenStudy (anonymous):

@Squickers I just told you what it is

OpenStudy (anonymous):

oh sorry didn't see that

OpenStudy (anonymous):

haha it is okay :)

OpenStudy (anonymous):

welcome to os :D

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

do you need anymore help???

OpenStudy (anonymous):

nah I figures it out

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\sqrt{x-2}=3\] for tristan

OpenStudy (anonymous):

\[\sqrt{x-2}=3\] for Hadley

OpenStudy (anonymous):

for tristan when you solve for x it gives you 11

OpenStudy (anonymous):

then if replace 11 for x and simplify it gives 3=-3

OpenStudy (anonymous):

meaning its extraneous

OpenStudy (anonymous):

If you do the same thing for hadley it gives you 11 again

OpenStudy (anonymous):

but when you replace x with 11 and simplify it gives you 3=3, not extraneous